OpenStudy (anonymous):

Simplify the expressions for questions 3–8. 3. 4^5/ 4^2 (1 point)

OpenStudy (raden):

use the property of exponent : a^m : a^n = a^(m-n)

OpenStudy (raden):

so, 4^5/4^2 = 4^(? - ?) = 4^ ?

OpenStudy (anonymous):

thank you soo much

OpenStudy (anonymous):

4. 120m^16/45m^8 (1 point)

OpenStudy (raden):

120/45 = ... (simplify) m^16/m^8 = ... (use the rule like no. 3 then multiply of both